OpenStudy (anonymous):

Perform the indicated operations. Subtract (3x - 30) from the sum of (-2x + 1) and (9x - 9) a. 8x + 40 b. 10x - 38 c. -4x - 22 d. 4x + 22

OpenStudy (oaktree):

4x+22, or D.

OpenStudy (oaktree):

-2x+1+9x-9 = 7x-8 7x-8 - (3x-30) = 7x-8-3x+30=4x+22.